OpenStudy (anonymous):

Give an example of two cars that have the same speed but different velocities.

OpenStudy (anonymous):

Is it like 10m/s and -10m/s

OpenStudy (across):

A car going in a straight line at 70 mph and a car running in circles at 70 mph.

OpenStudy (anonymous):

Okay so what would be the velocity of the one going in circles?

OpenStudy (across):

Let's make it simpler: A car going north at 70 mph and a car going south at 70 mph.

OpenStudy (anonymous):

Okay, I get it but I'm still curious what the velocity for the circle one would be? If you know

OpenStudy (across):

You can parametrize its position like this:\[r(t)=70\cos t+70\sin t,\]and thus its velocity will be:\[v(t)=-70sin t+70\cos t.\]

OpenStudy (anonymous):

Okkkkk ty

OpenStudy (across):

I forgot to multiply the terms by the unit vectors \(\hat i\) and \(\hat j\), but you get the idea.